# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ Acknowledgments
$\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}=\mathds{E}\|\mathcal{V}\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}\leq\|\mathcal{V}\|\mathds{E}\|\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=O(\mu^{2\gamma})$ (596)
Then similar to (E)–(327), we obtain
$\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}\leq O(\mu^{\gamma}),\quad\quad|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|\leq O(\mu^{1.5\gamma})$ (597)
@@
$\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}\leq O(\mu^{1+\eta}),\quad\quad|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|\leq O(\mu^{1.5(1+\eta))})$ (598)
Appendix J Additional simulation results
@@
In this section, we show the results associated with WideResNet-28-10 and DenseNet-121 on CIFAR10 and CIFAR100. We simulate these two neural networks with batch size $B=256$. In Figures 6 and 7, we illustrate the optimization performance and flatness separately. We further show the test accuracy of the three methods in Table III, from which we observe again that diffusion consistently outperforms the other two methods. Similar to ResNet-18, this phenomenon can be interpreted by the trade-off between flatness and optimization. For the comparison between decentralized and centralized methods, the flatness plays an important role since it can be observed from Figure 6 and 7 that decentralized training methods enable flatter models than the centralized approach without significantly losing optimization performance. However, as for the comparison between diffusion and consensus, the optimization performance values more as their flatness is very close. Specifically, we observe from Figure 6 that diffusion has smaller training loss than consensus. All in all, diffusion achieves a favorable balance between optimization and generalization among the three methods.
@@
TABLE III: Test accuracy of distributed algorithms on CIFAR10 and CIFAR100 with WideResNet-28-10 and Densenet-121.
@@
Figure 6: The evolution of training loss with WideResNet-28-10 and DenseNet-121.
@@
Figure 7: The flatness of WideResNet-28-10 and DenseNet-121.

### v2

@@ Acknowledgments
$\lim\limits_{n\to\infty}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}-\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}=\mathds{E}\|\mathcal{V}\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}\leq\|\mathcal{V}^{2}\|\mathds{E}\|\mathcal{V}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=\|\mathcal{V}^{2}\|\mathds{E}(\|\bar{\boldsymbol{z}}_{n}\|^{2}+\|\check{\boldsymbol{z}}_{n}\|^{2})\leq O(\mu^{2\gamma})$ (632)
Then similar to (D)–(345), we obtain
$\lim_{n\to\infty}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}\leq O(\mu^{\gamma}),\quad\quad\lim_{n\to\infty}|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|\leq O(\mu^{1.5\gamma})$ (633)
@@
$\lim\limits_{n\to\infty}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}\leq O(\mu^{1+\eta}),\quad\quad\lim\limits_{n\to\infty}|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|\leq O(\mu^{1.5(1+\eta))})$ (634)
Appendix I Additional simulation results
In this section, we detail the experimental setup for our study. The image classification tasks are performed on the CIFAR-10 and CIFAR-100 datasets. For data preprocessing, we adopt the standard data augmentation scheme, and normalize the data using channel-specific means and standard deviations, as outlined in [55, 56]. In decentralized training, the full dataset is randomly split into 16 subsets, each assigned to one local model, which only observes its respective subset during training. Centralized training follows the conventional single-agent learning approach. All decentralized experiments are simulated on a single Tesla V100 32GB GPU. For centralized experiments, due to CUDA memory limitations, the models are distributed across four Tesla V100 GPUs.
We evaluate the three distributed methods across various neural network architectures, graph structures, and batch sizes to ensure consistent and reliable results. The chosen neural network architectures include ResNet-18, WideResNet-28-10, and DenseNet-121. For the graph structures in the multi-agent system, simulations are conducted on two distinct topologies: (1) A random graph generated using the Metropolis rule [6] (Figure 3(a)). (2) A ring graph structure (Figure 3(b)). Both topologies consist of 16 nodes, with each node corresponding to a local model. ResNet-18 is evaluated on both random and ring graphs, while WideResNet-28-10 and DenseNet-121 are simulated only on the random graph. Regarding local batch sizes, we consider three configurations: 128, 256, and 512. In centralized training, the global batch size is scaled by a factor of $K$ (the number of nodes) relative to the local batch size. For example, when the local batch is 128 in decentralized methods, the corresponding global batch size for centralized training is $128\times 16=2048$. Additionally, ResNet-18 is evaluated on all of the three local batch size configurations, and WideResNet-28-10 and DenseNet-121 are simulated with the local batch 256.
For all experiments, the vanilla SGD optimizer is used to align with our theoretical findings, with the momentum parameter set to 0. We employ a piecewise learning rate schedule, starting with an initial learning rate of 0.2, which is reduced by a factor of 10 at $50\%$ and $75\%$ of the total epochs. The weight decay parameter is set to 0 for CIFAR-10 experiments and $41e-$ 4 for CIFAR-100 experiments. The number of training epochs varies across different settings, and they can be found from the x-axis of the plots depicting the evolution of training risk.
We now show the results associated with WideResNet-28-10 and DenseNet-121 on CIFAR10 and CIFAR100. We simulate these two neural networks with batch size $B=256$:
@@ Appendix I Additional simulation results
We first show the results corresponding to the scenario of training from scratch. We illustrate the optimization performance and flatness separately in Figure 8 and 9. We further show the test accuracy of the three methods in Table V, from which we observe again that diffusion consistently outperforms the other two methods. Similar to ResNet-18, this phenomenon can be interpreted by the trade-off between flatness and optimization. For the comparison between decentralized and centralized methods, the flatness plays an important role since it can be observed from Figure 8 and 9 that decentralized training methods enable flatter models than the centralized approach without significantly losing optimization performance. However, as for the comparison between diffusion and consensus, the optimization performance values more as their flatness is very close. Specifically, we observe from Figure 8 that diffusion has smaller training risk than consensus. All in all, in the scenario of training from scratch, diffusion achieves a favorable balance between optimization and generalization so that provides the best accuracy among the three methods.
@@
TABLE V: Test accuracy of distributed algorithms trained from scratch on CIFAR10 and CIFAR100 with WideResNet-28-10 and Densenet-121.
@@
Figure 8: The evolution of training risk with WideResNet-28-10 and DenseNet-121 when training from scratch.
@@
Figure 9: Flatness visualization of models trained from scratch with WideResNet-28-10 and DenseNet-121.
We next present the results of the scenario where decentralized algorithms are initialized at pretrained models. The optimization performance and flatness are depicted in Figure 10 and 11. Additionally, the test accuracy is shown in Table VI. Similar to the results of ResNet-18, decentralized methods converge to flatter minima than the centralized method without losing significant optimization performance, enabling them achieving higher test accuracy. Furthermore, when we delve more into the comparison between consensus and diffusion, consensus exhibits better test accuracy by converging to flatter minima while maintaining comparable optimization performance. In summary, in the scenario of pretrained initialization, consensus strikes the most favorable balance between optimization and generalization among the three methods.
TABLE VI: Test accuracy of distributed algorithms using pretrained initialization on CIFAR10 and CIFAR100 with WideResNet-28-10 and Densenet-121.
Dataset Architecture Centralized Consensus Diffusion
CIFAR10 WideResNet-28-10 $89.76\pm 0.34\%$ $\textbf{91.27}\pm 0.26\%$ $90.96\pm 0.41\%$
DenseNet-121 $89.23\pm 0.12\%$ $\textbf{90.42}\pm 0.26\%$ $90.09\pm 0.17\%$
CIFAR100 WideResNet-28-10 $67.61\pm 0.20\%$ $\textbf{71.30}\pm 0.22\%$ $69.82\pm 0.12\%$
DenseNet-121 $64.28\pm 0.12\%$ $\textbf{67.45}\pm 0.20\%$ $66.50\pm 0.16\%$
(a) WideResNet-28-10, CIFAR10
(b) WideResNet-28-10, CIFAR100
(c) DenseNet-121, CIFAR10
(d) DenseNet-121, CIFAR100
Figure 10: The evolution of training risk with WideResNet-28-10 and DenseNet-121 using pretrained initialization.
(a) WideResNet-28-10, CIFAR10
(b) WideResNet-28-10, CIFAR100
(c) DenseNet-121, CIFAR10
(d) DenseNet-121, CIFAR100
Figure 11: Flatness visualization of models using pretrained initialization with WideResNet-28-10 and DenseNet-121.